Reduce the order of the following differential equation and solve 23 %> + 22 – 32 = 3,3 > 0.

Please show detailed steps and in a clear writing. Thanks
Answer #1

2x+-3 z= 2 3 Let z then z=y and z= 2xy+y- 2 3 (22D°+xb-3) 2. 3 Put =e then xi D(D- fo=u3 emd D D -34 (2A(A-1)+D-3)7 =34 2 e 2-D-3 e 34 -34 2.943-2 9x3 yo4) = C22+ + 시2 C2x + 32+ -3+1 -3+1 Zazeiby 12 + -2 Ca -2 22 11